OpenStudy (anonymous):

write eq for line passing through point (4,12) and parallel to line x+2y=2

OpenStudy (anonymous):

Slope of the line is -1/2 I hope you can proceed now

OpenStudy (anonymous):

Now use y-12=-1/2(x-4) just simplify it

OpenStudy (anonymous):

what does final eq look like

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

y=-1/2+8?

OpenStudy (anonymous):

?